OpenStudy (cedric):

Choose the correct simplification of the expression (−5a^5b^4c^3)0.

OpenStudy (anonymous):

Hint: Any number times 0 is 0

OpenStudy (anonymous):

is it (−5a^5b^4c^3) * 0 or (−5a^5b^4c^3)^0

OpenStudy (anonymous):

Also, Hint: any finite number to the power 0 is 1

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

Eg: 10000000 *0 =0 0 *0 = 0

OpenStudy (anonymous):

Eg : 10000000^0 = 1 1^0 = 1